Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In B configuration; able to support 32+ baseline board pairs (16 standard; 32+ case-by-case). Many testing reports from last week.
  • Key dates:
    • 07-10 Mar: New Worlds, New Horizons (Santa Fe) - report from participants at Science Tea
    • 07 May - 15 Jun: 4-band science observing
  • Staff Travel:
  • Visitors:
    • 05 Jan - 06 Apr: John Cannon sabbatical at DSOC
    • 28 Feb - 28 Mar: Bob Sault

Hardware status and updates (Langley, Perley, Rupen)

Software status and updates (Butler)

  • Links:
  • New Executor version; OPT update coming (new hardware in the building; available for production ~2-3 weeks). CBE issues - memory/missing scans leading to a substantial fraction of "down-time" (noted in archive log).
  • CASA Update (new functionality available/issues/enhancements):Notes (Miriam): Code freeze today for 3.2; delivery 15 April; final meeting for 3.3 priorities is this Wednesday.
  • AIPS Update (new functionality available/issues/enhancements): Notes (Eric on travel through the 19th); switched power heuristics essentially converged (Dhawan/Perley/Owen).
  • SDM Update: Notes:

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon
    • 1030-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue
    • 0900-1700: Testing/software development time
    • 1800-0630: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu, Fri:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

  • Strong GC LST pressure (fixed data); reviewing possible science observing extensions on Tues and/or Fridays pending testing demands.

Issues Affecting Science Observing:

  • Correlator: fshift sub-band anti-aliasing causing phantom spectral lines when using recirculation for narrow bands (only worry if Michael has to prep your file); Bob's report from above.
  • *Correlator*: Missing/mislabeled scans in data sets (small percentage but recurring). . McMullin/Sowinski
  • Correlator: time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags);Report (McMullin).
  • Array/Construction: Frequency offset tests (Dhawan/Rupen): Recent tests.
  • Correlator: Delay rocking across bands. Brent investigating; tests to come.
  • Correlator: Subset of BlBs misbehaving (only affects hardware stacking modes). Currently restricted to 16 working Baseline boards; several RSRO projects are pending.
  • Antenna/High Freq: Reference pointing accuracy (C, X, Ku). Believed to be the result of a set of poor collimation settings on several antennas; updated.
  • Array: Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).

Commissioning Groups

Testing This week

  • Correlator: Rupen: Subarray testing; CBE merge??
  • Solar: Perley/Bastian: Solar observing tests (20dB testing); Software control of attenuator now available.
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Spectral line: Momjian: Zeeman test follow-up

RSRO/ECSO Reports

RSRO B configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011BRSROStatus

Individual reports (updates)

Misc:

Demo Science Status

AOB

  • Looking at a public web-based observing log (searchable) (Butler/McMullin).

Misc Actions

  • [Frail/Butler]: Enable restrictions based on proximity to sun; review OST weather trending.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised. - Sault visit * Low Priority [TBD]: Phase (and amp) wobbles; long history.fshift investigation.
-- JosephMcMullin - 2011-03-03
Topic revision: r4 - 2011-03-15, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback